# Clinically inspired enhance Explainability and Interpretability of an AI-Tool for BCC diagnosis based on expert annotation

### v2

@@ Abstract
Background and Objective: An AI tool has been developed to provide interpretable support for the diagnosis of BCC by teledermatology, thus speeding up referrals and optimizing resource utilization. Dermatologists base their BCC diagnosis on the presence or absence of specific BCC dermoscopic features. Our clinically inspired explainability approach relies on two key pieces of evidence: the BCC dermoscopic patterns identified in the image by the AI classification system and the image regions where the AI tool focuses, which correspond to the presence of BCC patterns.
Methods: A dataset of 1559 dermoscopic images was annotated by several dermatologists. MobileNet-V2 was employed for BCC/non-BCC classification and BCC pattern detection, developing a clinically-inspired XAI approach. The model was trained using stratified cross-validation and data augmentation to address class imbalance. Visual XAI was implemented to validate the clinically-inspired XAI detection. Since there is no established ground truth for BCC dermoscopic features, a standard reference was inferred from the diagnoses of four dermatologists using an algorithm based on expectation maximization (EM).
Results: The results demonstrate significant improvements in classification accuracy and interpretability, positioning this approach as a valuable tool for early BCC detection and referral to dermatologists. The BCC/non-BCC classification achieved an accuracy rate of 90%. For clinically-inspired explainable AI (XAI) results, the detection of BCC patterns useful to clinicians reaches 99% accuracy. As for the Clinically-inspired Visual XAI results, the mean of the Grad-CAM normalized value within the manually segmented clinical features is 0.57, while outside this region it is 0.16. This indicates that the model struggles to accurately identify the regions of the BCC patterns. These results demonstrate the ability of the AI tool to provide a useful explanation.
Conclusions: Based on these results, two significant clinical achievements have been attended: (1) A tool with dual explanation capabilities rooted in clinical practice has been developed, which improves reliability and adoption by dermatologists. (2) This advancement allows integration into the clinical field, which could change the current protocol in primary care centers to accelerate treatment for patients with BCC lesions in various areas of Andalusia.
@@
Basal Cell Carcinoma , Deep Learning , Skin Cancer , Telemedicine , XAI
journal: Artificial Intelligence in Medicine
[…]
[1]organization=University of Seville, city=Seville, country=Spain
[…]
[2]organization=Hospital Universitario Virgen de Macarena, city=Seville, country=Spain
[…]
[3]organization=Hospital Quirón Salud Infanta Luisa y Sagrado Corazón, city=Seville, country=Spain
@@ 1 Introduction
Worldwide, doctors diagnose skin cancer most commonly, with Melanoma, Basal Cell Carcinoma (BCC), and Squamous Cell Carcinoma (SCC) as the most prevalent types. BCC accounts for approximately $75$% of all skin cancers, making it the most frequent form. It has well-established clinical criteria for its diagnosis, yet there is significant variability in the presence of these clinical criteria between cases [1, 2, 3, 4].
@@
The development of public databases has significantly increased the number of published articles related to automatic detection of skin diseases in recent years [5, 6, 7, 8]. Although these databases are accessible and comprehensive for everyone, the clinical criteria used for diagnosis are not readily available. To develop a tool that is useful from a medical perspective, it is crucial to provide not only classification metrics, but also a detailed diagnosis explaining the detected clinical features, offering a more comprehensive result. In this context, Serrano et al. [9] created a clinically inspired skin lesion classification tool by detecting dermoscopic criteria for BCC.
@@
XAI techniques such as Grad-CAM [10] play a crucial role in improving the transparency and interpretability of CNNs in any kind of medical diagnosis. By offering visual explanations of the decision-making process, these methodologies allow medical professionals to gain insight into the predictive mechanisms of AI models, thus fostering trust and acceptance in AI-driven diagnoses. The integration of XAI with dermatological applications not only enhances the model’s interpretability but also aids in the early and accurate diagnosis of skin conditions, potentially leading to better patient outcomes. In van der Valden et al. survey [11] there are several works focused on the diagnosis of skin cancer that use different XAI techniques to see how their models behave are explained.
Barata et al. have developed a Computer-Aided Diagnosis (CAD) system based on deep learning for the hierarchical diagnosis of skin lesions, using ISIC 2017 and 2018 datasets. Their model uses attention modules to improve explainability by highlighting important features in dermoscopy images. This approach improves the transparency and interpretability of the diagnostic process, making the system more reliable and suitable for clinical use [12].
Chanda et al. developed a multimodal XAI system to support dermatologists in diagnosing melanoma. Using the HAM10000 dataset, which comprises over 10,000 dermoscopic images, they created a model that provides interpretable, text and region-based explanations alongside its diagnostic predictions. A three-phase reader study involving 116 international dermatologists demonstrated that the XAI system significantly increased clinicians’ trust and confidence in their diagnosis compared to a conventional AI system, although it did not significantly improve diagnostic accuracy [13].
Rezk et al. developed an interpretable skin cancer diagnosis model using clinical images to assist general practitioners in early detection and referral. Their approach incorporates a skin lesion taxonomy into a deep neural network to incrementally learn dermatological concepts. They used a dataset of 914 clinical images, supplemented by additional images from DermNet NZ to address class imbalances. To enhance transparency and trustworthiness in line with recent regulatory requirements, XAI techniques, such as Grad-CAM++, were employed to provide visual explanations of the model’s decisions. The proposed model significantly improved classification accuracy and interpretability, facilitating better understanding and detection of misdiagnoses [14].
However, in [12, 14] works, the use of XAI is employed to visualize the regions where the models have focused to extract features, providing a visual representation of the results returned by the classifier. Besides, [13] introduces a novel use of Grad-CAM to improve dermatologist trust by ensuring the AI model focuses on clinically relevant lesion areas. Grad-CAM explanations are combined with feature detection, aligning AI decisions with expert reasoning, significantly increasing user confidence without sacrificing diagnostic accuracy. Nevertheless, the key contribution of this work lies in a dual-layer explanation aimed at enhancing clinical diagnosis. Firstly, it offers a binary detection between BCC and non-BCC cases. Secondly, it provides a clinically valuable XAI tool for specialists, enabling them to verify whether its attention is being directed towards the correct BCC-related patterns.
@@ 2 Methodology
In modern healthcare, primary care physicians use teledermatology to receive high-quality diagnostic images remotely, allowing preliminary diagnoses of Basal Cell Carcinoma (BCC) using established patterns [4, 15]. Suspected BCC cases are immediately referred to dermatology specialists, improving healthcare efficiency, reducing waiting times, and facilitating early intervention.
Teledermatology is crucial in areas with limited specialist access and its integration with AI tools promises to enhance diagnostic accuracy and efficiency. This work aims to develop an AI tool to assist in this process by providing a binary classification of BCC/non-BCC with interpretable results Fig. 1 illustrates the current and proposed diagnostic workflows. In the proposed workflow, the AI tool not only delivers a BCC/non-BCC diagnosis but also provides a layer of interpretability, guiding the dermatologist to focus on the dermoscopic features most relevant to basal cell carcinoma. This dual-layer explanation ensures that the specialist can validate the AI’s predictions by confirming that the model is emphasising the correct clinical patterns, thus reducing the chances of diagnostic errors and increasing trust in AI-assisted teledermatology.
Figure 1: Illustration showing an example of the current workflow of BCC diagnosis by teledermatology in orange and the proposed workflow in green.
@@ 2.1 Database
The Dermatology Unit of the “Hospital Universitario Virgen Macarena” provided the entire database, which consists of dermoscopic images sent from 60 primary care centers during 2 years. The dataset comprises 1559 dermoscopic images divided into 3 subsets. Four dermatologists provided different types of annotation according to the subsets. Specifically:
  • 1.
    The first subset consists of 1089 dermoscopic images that were annotated following a multilabel codification explained in Sect. 2.1.1. To consolidate the multi-rater labeling, an Expectation-Maximization (EM) algorithm was employed, generating an inferred Standard Reference (SR), which will be explained in Sect. 2.1.2. Subsequently, a binary classification scheme was established to classify into BCC and non-BCC lesions.
  • 2.
@@
  • 3.
    The final subset was gathered from the ISIC Archive [8], utilizing 136 non-BCC images, predominantly consisting of nevus lesions.
@@ 2.1.1 Label codification
Each image may contain multiple dermoscopic patterns. Therefore, we used a one-hot coding scheme to encode the labels during image annotation and subsequently process the dermatologists’ annotations. Each image label is a binary word and each BCC dermoscopic pattern is a digit, where $1$ means presence and 00 means absence. The seven patterns (Fig. 2) that can appear in a BCC lesion are[4]: Pigment Network (PN) (negative criterion), Ulceration (U), Ovoid Nests (ON), Multiglobules (MG), Maple Leaf-like (ML), Spoke Wheel (SW), Arborizing Telangiectasia (AT). Thus, each label is a vector of dimensions $[1\textrm{x}7]$ using a binary codification where a value of 1 means BCC lesion and a value of 0 means non-BCC lesion. In Table 1 there are some examples of this process.
@@
2.1.2 Standard Reference (SR) Inferring
The accepted ground truth (GT) for BCC diagnosis is biopsy. However, there is no established GT for BCC dermoscopic patterns, which are subjectively assessed by dermatologists. Several studies have reported a low kappa coefficient when measuring inter-dermatologist agreement in determining the different dermoscopic patterns present in a lesion [16, 15]. Therefore, an adequate SR inferred from the consensus of several dermatologists is required. We implemented an Expectation-Maximization (EM) based algorithm [17] to derive a single SR for model training from multiple specialist labels. This algorithm consolidates multilabel annotations from different dermatologists and generates a single inferred SR that encapsulates the collective expertise of the raters. Silva et al. [18] used this algorithm to infer the SR from BCC pattern annotations and demonstrated that integrating this diverse expertise mitigates the subjectivity inherent in diagnosing the BCC pattern and improves the reliability and robustness of the classification model.
@@ 2.1.2 Standard Reference (SR) Inferring
BCC Non-BCC Pigment Network Ulceration Ovoid Nests Multiglobules Maple Leaf-like Spoke Wheel Arborizing Telangiectasia
@@
Table 2 summarises the distribution of labels in the database. As can be seen in this table, the database has a significant class imbalance, with SW and MG patterns under-represented. Several techniques have been used to address this problem, see Sect. 3.1.
2.2 Clinically-inspired XAI: BCC diagnosis with additional label information
This section refers to the AI model positioned at the bottom of Novel telemedicine diagnosis in Fig. 1 (Pattern detection AI). From a clinical point of view, achieving 100% accuracy in the detection of dermoscopic BCC patterns is not crucial to a correct diagnosis of BCC. A dermatologist only needs to detect one of the aforementioned BCC patterns to decide whether the lesion is BCC. Therefore, a useful XAI should present the same criterion. The proposed tool presents the following results: if no pattern is detected, it predicts no BCC; if the PN pattern is detected, it serves as a negative criterion, obtaining a prediction of no BCC; and if any other BCC pattern is detected, the prediction is positive for BCC.
The proposed AI tool is based on the MobileNet-V2 model with three classifier layers. The model was coupled with a three-step optimization and training strategy. The first two steps focused on the detection of BCC/non-BCC, while the third step focused on pattern classification. In the first step, ImageNet transfer learning was applied and classifier weights were trained. In the second step, fine-tuning was applied to the last three blocks of the feature extractor, and the classifier was tuned with a lower learning rate (LR) and an increased number of epochs. In the third step, once the column extracts features from the same region as the specialist in the binary task, the classifier was retrained for pattern detection with a very low LR.

### v3

@@ Abstract
Purpose: Basal cell carcinoma (BCC) accounts for 75% of all skin cancers. Currently, all major public hospitals in Spain have a dermatology care protocol that includes teledermatology. This has created an overload for hospital dermatologists, which could be alleviated with an AI tool for prioritization. Several AI systems have been proposed for this purpose, but the lack of transparent diagnostic explanations limits their clinical acceptance and implementation. This fact motivates the present study, which aims to develop an AI tool focused on detecting BCC from dermoscopic images incorporating dermatologist diagnostic criteria to enhance reliability. Specifically, the BCC diagnostic criterium is that a lesion is not considered BCC if it exhibits pigment network pattern, and that a lesion is considered BCC if it exhibits at least one of these BCC patterns: ulceration, ovoid nest, multi globules, maple-leaf, spoke wheel, arborizing telangiectasia.
Methods: We analyzed 1,559 dermoscopic images collected from 60 primary care centers in Andalusia. Four dermatologists annotated the images as exhibiting or not each of the seven possible BCC patterns. As there is no established Ground Truth to determine the BCC patterns present in a lesion, we propose an Expectation-Maximization consensus algorithm to consolidate the multi-rater annotations into a unified standard reference (SR). As an additional novelty, the system incorporates the symbolic reasoning of dermatologists, who base their diagnoses on BCC patterns shown in lesions. To this end, a multitask learning (MTL) system based on MobileNet-V2 was designed. This system can rapidly triage BCC and non-BCC lesions while providing clinical information justifying this classification. This system also provides GradCAM-based maps to dermatologists to improve its reliability and confidence.
Results: Three evaluations were performed on the AI system. First, a performance analysis was conducted to evaluate the AI tool’s ability to classify lesions as BCC or non-BCC. In this analysis, the model achieved 90% accuracy (precision=0.90, recall=0.89). The second evaluation analyzed whether the detected patterns agreed with dermoscopic criteria. Notably, at least one clinically relevant BCC pattern was correctly identified in 99% of BCC-positive cases, and the pigment-network negative criterion was met in 95% of non-BCC cases. A comparison of the GRADCAM maps with the dermatologist’s manual delineation demonstrated strong colocalization with dermatologist-segmented regions (mean foreground density 0.57 vs. background 0.16), confirming alignment of the visual focus of experts.
Conclusions: This work introduces the first clinically validated dual-explanation AI system that combines high-accuracy BCC detection with transparent, pattern-based explanations. This approach closes the critical gap between AI performance and clinical trust in teledermatology, positioning the system for immediate deployment in primary care. Future work will focus on determining the extent to which this dual explanation system improves dermatologists’ confidence.
@@
Basal Cell Carcinoma, Deep Learning, XAI, Dermoscopy, Telemedicine, Computer-Aided Diagnosis
@@ 1 Introduction
Worldwide, doctors diagnose skin cancer most commonly, with Melanoma, Basal Cell Carcinoma (BCC), and Squamous Cell Carcinoma (SCC) as the most prevalent types. BCC accounts for approximately 7575% of all skin cancers, making it the most frequent form. It has well-established clinical criteria for its diagnosis, yet there is significant variability in the presence of these clinical criteria between cases [aad2024, skincancerfoundation2024, cikazynska2018trends, peris2019diagnosis_BCCPatterns].
@@
The development of public databases has significantly increased the number of published articles related to automatic detection of skin diseases in recent years [Codella2017, combalia2019bcn20000, HAM10000DataBase, isicarchive2024]. Although these databases are accessible and comprehensive for everyone, the clinical criteria used for diagnosis are not readily available. To develop a tool that is useful from a medical perspective, it is crucial to provide not only classification metrics, but also a detailed diagnosis explaining the detected clinical features, offering a more comprehensive result. In this context, Serrano et al. [serrano2022] created a clinically inspired skin lesion classification tool by detecting dermoscopic criteria for BCC.
Current meta-analyses indicate that while AI algorithms significantly outperform general practitioners in skin cancer classification, achieving a sensitivity of 92.5% versus 64.6% for generalists, their performance remains comparable to that of expert dermatologists, suggesting that the optimal clinical role for AI is as a collaborative support tool rather than a standalone replacement [salinas2024]. To facilitate this integration, recent research has expanded the scope of AI beyond simple lesion classification. For instance, Liu et al. [liu2024] demonstrated the utility of AI in early screening by predicting skin cancer risk directly from 2D facial images, identifying complex non-linear facial endophenotypes that outperformed traditional risk factor assessments (c-index 0.72 vs. 0.59). Simultaneously, to address the “black box” limitations in lesion diagnosis, studies have begun to explicitly encode clinical reasoning into neural architectures. Wang et al. [wang2025] introduced a framework utilizing Clinical Knowledge Graphs to model the directed relationships between dermoscopic attributes of the 7-point checklist, thereby mimicking the structured decision-making process of a clinician.
@@
XAI techniques such as Grad-CAM [Selvaraju_2019_GradCAM] play a crucial role in improving the transparency and interpretability of CNNs in medical diagnosis. By offering visual explanations of the decision-making process, these methodologies allow medical professionals to gain insight into the predictive mechanisms of AI models, thus fostering trust and acceptance in AI-driven diagnoses.
Recent works in dermatological AI have employed various XAI approaches: attention-based mechanisms to highlight diagnostically relevant features [Barata_RNN_XAI], multimodal explanations combining text and visual information to increase clinician confidence [Chanda2024], and gradient-based visualization methods like Grad-CAM++ to provide post-hoc interpretability and facilitate detection of misdiagnoses [Rezk2023]. Van der Valden et al. [survey_XAI] provide a comprehensive survey of XAI techniques applied to skin cancer diagnosis. However, these works primarily focus on the explainability of the regions where models focus their attention, providing a post-hoc interpretation of the features extracted by a trained classifier. Thus, clinician trust increases when AI activation overlaps with clinically relevant regions.
In contrast, the key contribution of this work lies in a Multi-Task Learning (MTL) framework that integrates clinical reasoning directly into the model architecture. This dual-layer structure not only performs binary classification of BCC versus non-BCC lesions but also provides pattern-level interpretability grounded in dermoscopic features. By jointly learning these tasks, the system enables a clinically meaningful form of XAI, allowing dermatologists to validate the prediction rationale through pattern-specific attention analysis—thus enhancing the reliability, transparency, and usability of the tool in telemedicine workflows. The practical consequence of integrating clinical reasoning into the model architecture is that the proposed XAI model offers two clinical explanations. First, it provides an explanation of the BCC patterns encountered in the image. Second, it provides a GRADCAM map correlated with the location of these BCC patterns.
This research is driven by the need to improve clinical workflows in healthcare systems based on teledermatology. In modern healthcare, primary care physicians use teledermatology to receive high-quality diagnostic images remotely, allowing preliminary diagnoses of Basal Cell Carcinoma (BCC) using established patterns [peris2019diagnosis_BCCPatterns, peris2002interobserver_BCCPatterns3]. Teledermatology is crucial in areas with limited specialist access and its integration with AI tools promises to enhance diagnostic accuracy and efficiency. Suspected BCC cases are immediately referred to dermatology specialists, improving healthcare efficiency, reducing waiting times, and facilitating early intervention. This work aims to develop an AI tool to assist in this process by providing a binary classification of BCC/non-BCC with interpretable results.
@@ 2 Methodology
The proposed methodology to improve the teledermatology workflow is summarized in Fig. 1, which illustrates both the current and proposed diagnostic workflows. In the proposed workflow, the AI tool not only delivers a BCC/non-BCC diagnosis but also provides a layer of interpretability, guiding the dermatologist to focus on the dermoscopic features most relevant to basal cell carcinoma. This dual-layer explanation ensures that the specialist can validate the AI’s predictions by confirming that the model is emphasising the correct clinical patterns, thus reducing the chances of diagnostic errors and increasing trust in AI-assisted teledermatology.
Figure 1: Comparison between the current and proposed workflows for basal cell carcinoma (BCC) diagnosis via teledermatology. The upper panel (orange) illustrates the standard approach, in which the dermatologist performs manual pattern recognition and binary diagnosis. The lower panel (green) depicts the proposed multi-task learning (MTL) system, where a shared AI backbone simultaneously performs BCC/non-BCC classification and BCC pattern detection, supported by clinically-inspired visual XAI. This novel framework aims to increase automation, interpretability, and clinical alignment in remote diagnosis.
@@ 2.1 Database
The Dermatology Unit of the “Hospital Universitario Virgen Macarena” provided the entire database, which consists of dermoscopic images sent from 60 primary care centers during the period 2022–2023. The majority of images used in this study were collected between July and November 2023, coinciding with the deployment of the teledermatology system across these primary care centers. The dataset comprises 1559 dermoscopic images divided into 3 subsets. Four dermatologists provided different types of annotation according to the subsets. Specifically:
  • The first subset consists of 1089 dermoscopic images that were annotated following a multilabel codification explained in Sect. 2.1.1. To consolidate the multi-rater labeling, an Expectation-Maximization (EM) algorithm was employed, generating an inferred Standard Reference (SR), which will be explained in Sect. 2.2. Subsequently, a binary classification scheme was established to classify into BCC and non-BCC lesions.
@@
  • The final subset was gathered from the ISIC Archive [isicarchive2024], utilizing 136 non-BCC images, predominantly consisting of nevus lesions.
Images were acquired using different dermoscopic devices deployed across the primary care network, including portable polarized digital dermatoscopes (DermLite II Multispectral, 3Gen) with smartphone adapters [AGSSurSevilla2025, hospitalmacarena2019], and professional videodermatoscopy systems (FotoFinder Medicam 1000 with D-Scope III) at hospital level [aetsa2006telederma]. This device heterogeneity reflects real-world teledermatology conditions. Regarding image preprocessing, no specific image processing was applied; only images with insufficient quality due to motion blur, poor focus, or inadequate exposure were excluded. The inclusion criterion was adequate visualization of the lesion for clinical assessment. Patient demographic data (age, sex, skin type) were not available due to anonymization protocols required by the ethics committee.
@@ 2.1.1 Label codification
Each image may contain multiple dermoscopic patterns. Therefore, we used a one-hot coding scheme to encode the labels during image annotation and subsequently process the dermatologists’ annotations. Each image label is a binary word and each BCC dermoscopic pattern is a digit, where $1$ means presence and 0 means absence. The seven patterns (Fig. 2) that can appear in a BCC lesion are [peris2019diagnosis_BCCPatterns]: Pigment Network (PN) (negative criterion), Ulceration (U), Ovoid Nests (ON), Multiglobules (MG), Maple Leaf-like (ML), Spoke Wheel (SW), Arborizing Telangiectasia (AT). Thus, each label is a vector of dimensions $1\textrm{x}7]$ using a binary codification where a value of 1 means BCC lesion and a value of 0 means pattern absence. In Table 1 there are some examples of this process. In Examples 1 and 2, the labels indicate the presence of BCC dermoscopic patterns. According to clinical criteria, this corresponds to the presence of BCC, which is encoded with the binary label 1. Example 3 indicates the presence of a pigment network, while example 4 indicates that no BCC pattern can be found in the lesion. Thus, the binary encoding must be 0 in both cases, indicating an absence of BCC.
@@
2.2 Standard Reference (SR) Inferring
The accepted ground truth (GT) for BCC diagnosis is biopsy. However, there is no established GT for BCC dermoscopic patterns, which are subjectively assessed by dermatologists. Several studies have reported a low kappa coefficient when measuring inter-dermatologist agreement in determining the different dermoscopic patterns present in a lesion [polesie2021interobserver, peris2002interobserver_BCCPatterns3]. Therefore, an adequate SR inferred from the consensus of several dermatologists is required. We implemented an Expectation-Maximization (EM) based algorithm [dawid1979ExpectationMaximization] to derive a single SR for model training from multiple specialist labels. This algorithm consolidates multilabel annotations from different dermatologists and generates a single inferred SR that encapsulates the collective expertise of the raters. Silva et al. [Invent] used this algorithm to infer the SR from BCC pattern annotations and demonstrated that integrating this diverse expertise mitigates the subjectivity inherent in diagnosing the BCC pattern and improves the reliability and robustness of the classification model.
@@ 2.2 Standard Reference (SR) Inferring
BCC Non-BCC PN U ON MG ML SW AT
@@
Table 2 summarises the distribution of labels in the database. As can be seen in this table, the database has a significant class imbalance, with SW and MG patterns under-represented. Several techniques have been used to address this problem, see Sect. 2.4.
2.3 Multi-task learning (MTL) System: BCC Diagnosis with XAI-guided Pattern Recognition
This section refers to the AI model positioned at the bottom of the Novel telemedicine diagnosis framework in Fig. 1 (Pattern detection AI). The system was developed under a multi-task learning (MTL) paradigm, in which a single backbone jointly supports two interrelated objectives: the binary classification of BCC versus non-BCC lesions, and the detection of clinically relevant BCC dermoscopic patterns that guide this decision. From a clinical point of view, achieving 100% accuracy in the detection of BCC patterns is not required for a reliable diagnosis: a dermatologist only needs to identify one indicative pattern to confirm the presence of BCC. Accordingly, a clinically aligned XAI system should follow the same principle. The proposed model adheres to this logic: if no pattern is detected, it predicts no BCC; if the PN pattern is detected, it acts as a negative criterion; and if any of the other BCC patterns are detected, the system issues a positive prediction for BCC. The MTL architecture is based on the MobileNet-V2 model with three classifier layers. The model was coupled with a three-step optimization and training strategy. The first two steps focused on the detection of BCC/non-BCC, while the third step focused on pattern classification. In the first step, ImageNet transfer learning was applied and classifier weights were trained. In the second step, fine-tuning was applied to the last three blocks of the feature extractor, and the classifier was tuned with a lower learning rate (LR) and an increased number of epochs. In the third step, once the column extracts features from the same region as the specialist in the binary task, the classifier was retrained for pattern detection with a very low LR.